40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 6713     Entry time: Wed May 30 01:35:15 2012
Author: yuta 
Type: Update 
Category: Green Locking 
Subject: aligned Y arm green beam 

[Jenne, Yuta]

We aligned the Y arm for IR (C1:LSC-TRY_OUT is now ~ 0.9), and aligned the green beam from the ETMY table. The Y arm green is now resonating in TEM00 mode, but we need some monitors (green trans or green refl) to maximize the coupling.

We noticed that the MC beam spot are oscillating at ~ 1 Hz, mostly in YAW.  This wasn't observable before the PMC realignment (elog #6708). We should find out why and fix it.

ELOG V3.1.3-